OpenStudy (anonymous):

An architect built a scale model of a shopping mall. On the model, a circular fountain is 20 inches tall and 22.5 inches in diameter. If the actual fountain is to be 8 feet tall, what is its diameter? Plz help makes no sense.

OpenStudy (anonymous):

You would set up a ratio equality to convert the diameter.

OpenStudy (anonymous):

How would it look when set up?

OpenStudy (anonymous):

First, convert the 8 feet to inches. Then the ratio of the actual height to the model height equals the actual diameter to the model diamter.

OpenStudy (anonymous):

You use a fraction to get the ratio.

OpenStudy (anonymous):

When it is set up, it will look like one fraction equaling another.

OpenStudy (anonymous):

How do you convert the 8 feet to inches?

OpenStudy (anonymous):

How many inches in one foot? Then, how many inches in 8 feet?

OpenStudy (anonymous):

12.. so 8 * 12?

OpenStudy (anonymous):

Yes, good so far. So, now set up the first fraction of actual height to model height.

OpenStudy (anonymous):

96/ 20?

OpenStudy (anonymous):

Yes. Now that fraction, which is also conceptually a ratio, will be equal to the ratio of the actual diamter to the model diameter.

OpenStudy (anonymous):

so 96/20 = 22.5/ x?

OpenStudy (anonymous):

We are solving for the actual diameter, so just call it "x" for now. You are close but off a bit. You have to actual on top (numerator) for both sides. On the right side, you have it in the denominator.

OpenStudy (anonymous):

So, just flip over the right side.

OpenStudy (anonymous):

96/ 20 = x / 22.5?

OpenStudy (anonymous):

Yes, very good. You are almost done. You only have to solve for "x" now.

OpenStudy (anonymous):

96 * 22.5 = 2025 / 20 x = 101.25?

OpenStudy (anonymous):

I'll re-write your previous equation where it was still on track:\[\frac{ 96 }{ 20 } = \frac{ x }{ 22.5 }\]gives:\[\frac{ 96 \times 22.5 }{ 20 } = x\]

OpenStudy (anonymous):

It's not 101.25

OpenStudy (anonymous):

ah, think I change a number or something.. x = 108

OpenStudy (anonymous):

You got it! Good job!

OpenStudy (anonymous):

:)! U sure its 108.. just asking cuz its not one of my choices I have 105 but not 108

OpenStudy (anonymous):

If you wrote out the problem correctly, then 108 is definitely the answer.

OpenStudy (anonymous):

Yeahh I copied and paste it. Um, that's weird.

OpenStudy (anonymous):

Sometimes the answer key is off because it could have a typo. That doesn't happen often, but it does happen.

OpenStudy (anonymous):

Wait I don't even have 105 its 10.5.. u sure that 108 is the diameter?

OpenStudy (anonymous):

108 inches.

OpenStudy (anonymous):

This problems a pain lol

OpenStudy (anonymous):

108 inches is 9 feet. Is one of your answers "9"?

OpenStudy (anonymous):

:) Yeah!

OpenStudy (anonymous):

Then you have "9" for your answer. Looks like the key is right after all, and so was my answer (both ways).

OpenStudy (anonymous):

Good luck to you in all of your studies and thanks for the recognition! @LilVeroDez13

OpenStudy (anonymous):

:) yes it was. thanks a lot!! would u mind helping me with one more? :)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Thanks! and ur welcome (:

OpenStudy (anonymous):

Just start a new thread.

OpenStudy (anonymous):

Ok :)